OpenStudy (anonymous):

is x^2y^3z/xy^2 equivanent to xz/y^-1

OpenStudy (phi):

do you know how to simplify fractions with exponents?

OpenStudy (anonymous):

not really

OpenStudy (campbell_st):

using index laws for division x^(2 -1)y^(3-2)z^1 gives xyz and \[xz/y^-1 = xz \times y = xyz\] they are equvalent

OpenStudy (campbell_st):

the key is \[1/y^{-1} = y\]